Sally brought three chocolate bars and a pack of gum and paid $1.75. Jake brought two chocolate bars and four packs of gum and paid $2.00 Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by stanbon(75887)

Sally brought three chocolate bars and a pack of gum and paid $1.75. Jake brought two chocolate bars and four packs of gum and paid $2.00 Find the cost of a chocolate bar and the cost of a pack of gum.
--------
Let cost of cb's be "x" and cost of pg's be "y".
EQUATIONS:
1st: 3x+y=1.75
2nd: 2x+4y=2.00
Multiply 1st by 4 to get
3rd: 12x+4y=7
Subtract 2nd from 3rd to get:
10x=5
x=1/2= 50 cents
Substitute into 2nd to solve for "y":
2(1/2)+4y=2.00
4y=1
y=1/4= 25 cents
Cheers,
Stan H.